OpenStudy (anonymous):

Solve for w=z^2 (complex transformation)

OpenStudy (amoodarya):

write z as x+iy w=z^2=

OpenStudy (amoodarya):

w=(x+iy)^2 =x^2+i^2y^2+2ixy w=(x^2-y^2) +i( 2xy)